the energy difference between the ground level and the first vibration level of oxygen gas is equal to 10^-29 x 3.1391 joules. Compare the ratios of the particles between two levels at 750 K and 2000 K , and calculate the wave number for two temperatures


1
Expert's answer
2021-02-02T15:34:20-0500

"W = 3.1391*10^{-29}"

"T_1 = 750K \space \space T_2 = 200K"

Oxygen "O_2" "M(O_2) = 32g\/mol"

W = "\frac{3}{2}\nu R \delta T+ \frac{hc}{\lambda}" "\to \lambda= \large\frac{hc}{W - \frac{3}{2}\nu R \delta T} = \frac{6.62*10^{-34}*3*10^8}{3.1391*10^{-29} -1.5*32*10^{-3}*8.31*1250} =" 832nm